OpenStudy (anonymous):

36 - 3X = 45 -5 -3 15 -2

OpenStudy (anonymous):

36 - 3x = 45 (subtract 36 from each side -3x = 45 - 36 -3x = 9 (now divide both sides by -3) -3x/-3 = 9/-3 x = - 3